# The Cognitive Foundations of Economic Exchange: A Modular Framework Grounded in Behavioral Evidence

### v1

@@ Abstract
A key challenge in multi-agent AI is modeling social cooperation under realistic behavioral constraints. Many foundational concepts in economics and ethics—such as ”trust” or ”morality”—are often defined informally, without operational criteria or cognitive grounding, which limits their testability and implementation in artificial agents. Drawing on converging empirical evidence from primate behavior, infant cognition, and economic anthropology, we propose a conceptual framework composed of three cognitively minimal mechanisms: individual recognition, reciprocal credence, and cost–return sensitivity. This framework reframes trust as a graded cognitive expectation, providing a simulateable basis for reciprocal exchange in artificial agents, and enabling the bottom-up emergence of scalable cooperation and institutional dynamics.
@@ 1 Introduction
In this paper, we argue that the foundations of economic exchange—and by extension, scalable cooperation—can be traced to three cognitively minimal, biologically grounded mechanisms: individual recognition, reciprocal credence, and cost–return sensitivity. These capacities are observed across human and nonhuman primates, and support sustained reciprocal exchange without requiring symbolic trust or formal enforcement.
@@
We offer a theoretical realignment: a cognitively grounded account of reciprocal exchange that clarifies how economic systems can emerge from social behavior rather than institutional design. Specifically, we:
@@
This work reframes economic modeling not as a top-down implementation of rules, but as a bottom-up emergence of institutions from cognitive interaction.
@@ 2.2 Reciprocal Behavior in Nonhuman Primates
Our framework builds on these findings by situating primate reciprocity within a cognitive-behavioral model. We argue that capacities such as individual recognition, memory of interaction history, and fairness sensitivity form a biologically grounded substrate upon which more complex human economic systems could emerge.
@@ 2.4 From Behavioral Anomalies to Cognitive Substrates
In contrast, our framework treats these preferences as emergent outcomes of specific cognitive-social mechanisms: individual recognition, memory of interaction, and cost–return evaluation. By shifting the focus from outcomes to underlying capacities, we aim to bridge behavioral economic observations with a biologically grounded account of exchange behavior.
@@ 3.1 The Myth of Barter and the Reciprocal Foundations of Exchange
3.2 The Three Core Mechanisms of Biological Reciprocity
@@
These three capacities are supported by converging evidence from nonhuman primates and human infants, suggesting they are both biologically plausible and developmentally early. Importantly, they are not merely descriptive—they are simulateable as modular components in multi-agent systems.
While primate studies have richly documented reciprocal behaviors—grooming, food sharing, coalition support—the cognitive prerequisites for sustaining such behavior are often left implicit. As (Schino & Aureli, 2010) note, reciprocity is frequently treated as a local response pattern, rather than a structured process involving memory and expectation. Yet reciprocal exchange, by definition, requires continuity. Without recognition or anticipation of future interaction, trust-like behavior cannot functionally emerge.
By defining these mechanisms as behavioral primitives, we offer a bottom-up model of economic emergence—one that grounds institutional systems not in rules, but in biologically plausible, cognitively minimal substrates of social interaction.
@@ 4 Theoretical Framework: Simulateable Primitives for Economic Exchange
Existing efforts in agent-based cooperation often rely on engineered incentives, fixed strategy spaces, or symbolic approximations of concepts like “trust.” What remains missing is a biologically grounded, simulateable account of the behavioral capacities that make reciprocal exchange possible.
We propose three minimal mechanisms—individual recognition, reciprocal credence, and cost–return sensitivity—as a compositional framework for simulating the emergence of economic exchange. These mechanisms are:
  • Cognitively minimal: each is supported by empirical data from nonhuman primates and human infants;
  • Behaviorally tractable: each has observable correlates in real-world cooperation;
  • Simulateable: each can be implemented in agent-based systems as a decision module or update function.
This framework allows social exchange to be modeled from the bottom up—without relying on predefined markets, institutions, or payoff matrices.
@@ 4.1 Individual Recognition
We define individual recognition as the ability to identify and distinguish specific agents across time, enabling the tracking of prior social interactions. This capacity underlies the first and most basic source of reciprocal credence: memory of positive experiences with known others. Without recognition, no information about past cooperation can be meaningfully retained or used to guide future behavior.
@@ 4.2 Reciprocal Credence
To address this gap, we introduce the concept of reciprocal credence: a cognitively grounded, graded expectation that another agent will respond to cooperation with cooperation. Reciprocal credence does not assume moral virtue or social obligation. Instead, it arises from informational cues—typically based on one of two elements: (1) past prosocial behavior directed toward the self or others, or (2) contextual signals that make cooperative return likely. It is not a feeling, but a function: a decision-weighting variable that governs the willingness to enter or continue reciprocal exchange.
Unlike categorical trust, reciprocal credence is inherently scalar: it reflects how likely an agent believes cooperation will be reciprocated, not whether it will be or not. Treating reciprocal credence as a scalar variable is essential to modeling cooperation as a dynamic, context-sensitive behavior.
We distinguish three primary sources of reciprocal credence:
@@
    Direct positive interaction history, where an individual has experienced prior cooperative behavior, prosocial engagement, or emotional bonding with another.
@@
    Reputational inference, where the individual observes or learns about others’ cooperative behavior.
@@
    Role-based expectation, where social roles or institutional norms lead to default expectations of reciprocity.
Each source varies in reliability, stability, and cognitive cost, but all contribute to the emergence of trust-like behavior without requiring moral commitment. Together, these sources provide the informational basis for reciprocal credence—a flexible, graded expectation that others will act in ways conducive to cooperation.
@@
It is also important to distinguish social trust from what might be better described as functional dependency. Contemporary systems such as ChatGPT or Google are often described as being “trusted,” when in reality they are simply relied upon due to their consistent output, broad adoption, and lack of failure. This perceived “trust” is better understood as a product of usability, network effects, and repeated exposure — not reciprocal social accountability.
A similar semantic conflation is common in financial discourse, where “trust in money” or “trust in banks” is frequently cited as foundational to economic stability. However, what is usually meant in such contexts is not interpersonal trust, but confidence in the continuity of institutional enforcement, liquidity guarantees, or the absence of systemic failure.
@@ 4.3 Cost–return sensitivity
We define cost–return sensitivity as the capacity to evaluate whether engaging in cooperative behavior yields a net benefit over time. Unlike categorical heuristics (e.g., always cooperate or always defect), this sensitivity enables agents to condition their participation on the payoff structure of ongoing interactions. It allows withdrawal from exploitative relationships and reinforcement of beneficial ones, serving as a regulatory mechanism for reciprocal dynamics.
This capacity need not involve explicit calculation; approximate assessments—such as tracking recent imbalances in giving and receiving—are sufficient to support adaptive behavior. For example, chimpanzees are less likely to groom or share with partners who have failed to reciprocate in the past, even if those costs are not immediate or symmetrical. Similarly, infants gradually reduce helping behavior toward agents who consistently fail to reciprocate attention or aid.
Such sensitivity ensures that reciprocal exchange remains robust under uncertainty, preventing exploitation while maintaining flexibility in dynamic social environments.
@@
5 Possible Implementation via Memory Design
While each proposed mechanism is tractable in isolation, integrating them into a unified architecture remains an open challenge—particularly due to current limitations in memory design for LLM-based agents. Most existing multi-agent frameworks lack persistent, structured memory, resulting in shallow social modeling and limited capacity for behavioral adaptation.
We propose that each behavioral primitive—individual recognition, reciprocal credence, and cost–return sensitivity—can be realized through targeted extensions to agent memory and retrieval mechanisms:
  • Individual recognition can be implemented using identity-specific memory slots or embedding-indexed retrieval systems, enabling agents to store and retrieve interaction histories with named partners. This supports continuity in agent behavior and the ability to condition future responses on partner identity.
  • Reciprocal credence may be represented as an internal latent variable or scalar annotation associated with each agent in memory (e.g., “Agent X: credence = 0.7”), and updated through lightweight heuristics (e.g., recency-weighted prosocial acts). This allows LLM agents to infer, represent, and update partner expectations without requiring explicit reward functions.
  • Cost–return sensitivity can be implemented through structured memory logs of past cooperative attempts and outcomes. Agents may track cumulative give–take asymmetries and adjust future investments accordingly. Prompts can surface this information as part of decision rationale (e.g., “Agent Y has taken without giving for three turns”).
These strategies enable biologically inspired reciprocity to be modeled using memory-centric extensions, rather than reinforcement learning or symbolic planning. By shifting the burden of social behavior onto structured memory design and prompt-level reasoning, this framework opens a tractable path toward more socially competent LLM agents.
@@ 6.1 Limitations of the Framework
Although we argue that modern exchange institutions build upon a biologically grounded behavioral substrate, we do not claim that such mechanisms alone are sufficient to explain the emergence of formal systems like currency, interest-bearing debt, or taxation. These likely require additional cultural, historical, and symbolic developments.
@@ 6.2 Theoretical Implications
To address this, we introduce reciprocal credence: a graded, biologically grounded expectation that cooperation will be returned. It emerges from recognition, interaction history, and cost–return sensitivity—not from moral obligation. In contrast, large-scale “trust” in systems like money or platforms is better understood as confidence in robustness, not social accountability. This reframing clarifies how reciprocal behavior scales without assuming institutional trust.
@@ 6.3.2 Simulating Institution Formation via Behavioral Primitives
With the rise of large language model (LLM)-based agents and embodied multi-agent simulators (e.g., Habitat, ThreeDWorld, AutoGen), it is now possible to explore whether agents endowed with these primitives can spontaneously develop systems of reciprocal exchange, role specialization, symbolic debt, or proto-monetary behaviors. Such experiments would help identify the minimal substrate for scalable cooperation and provide a generative framework for building socially competent artificial agents.

### v2

@@ Abstract
A key challenge in multi-agent AI is modeling social cooperation under realistic behavioral constraints. Many foundational concepts in economics and ethics—such as ”trust” or ”morality”—are often defined informally, without operational criteria or cognitive grounding, which limits their testability and implementation in artificial agents. Drawing on converging empirical evidence from primate behavior, infant cognition, and economic anthropology, we propose a conceptual framework composed of three cognitively minimal mechanisms: individual recognition, reciprocal credence, and cost–return sensitivity. This framework reframes trust as a graded cognitive expectation, providing a simulateable basis for reciprocal exchange in artificial agents, enabling the bottom-up emergence of scalable cooperation, proto-economic norms, and institutional structure.
@@ 1 Introduction
In this paper, we argue that the foundations of economic exchange—and by extension, scalable cooperation—can be traced to three cognitively minimal, behaviorally grounded mechanisms: individual recognition, reciprocal credence, and cost–return sensitivity. These capacities are observed across human and nonhuman primates, and support sustained reciprocal exchange without requiring symbolic trust or formal enforcement.
@@
We offer a theoretical realignment: a cognitively grounded account of reciprocal exchange that clarifies how economic systems can emerge from social behavior rather than institutional design. We argue that reciprocity is not just a component of cooperation, but the minimal cognitive substrate from which stable social structures emerge.
@@
Unlike prior models that treat cooperation heuristically or assume trust as exogenous, our framework offers a cognitively minimal and simulateable account of how such expectations emerge from social interaction. This work reframes economic modeling not as a top-down implementation of rules, but as a bottom-up emergence of institutions from cognitive interaction.
@@ 2.2 Reciprocal Behavior in Nonhuman Primates
Our framework builds on these findings by situating primate reciprocity within a cognitive-behavioral model. We argue that capacities such as individual recognition, memory of interaction history, and cost-return sensitivity form a behaviorally grounded substrate upon which more complex human economic systems could emerge.
@@ 2.4 From Behavioral Anomalies to Cognitive Substrates
In contrast, our framework treats these preferences as emergent outcomes of specific cognitive-social mechanisms: individual recognition, memory of interaction, and cost–return evaluation. By shifting the focus from outcomes to underlying capacities, we aim to bridge behavioral economic observations with a behaviorally grounded account of exchange behavior.
@@ 3.1 The Myth of Barter and the Reciprocal Foundations of Exchange
3.2 The Three Behavioral Mechanisms Enabling Reciprocity
@@
These mechanisms are supported by converging evidence from nonhuman primates and human infants, suggesting they are biologically plausible, developmentally early, and cognitively minimal. Each is also behaviorally tractable and simulateable in multi-agent systems—whether as memory modules, value estimators, or policy regulators.
While primate studies have richly documented behaviors like grooming, food sharing, and coalition support, the cognitive prerequisites for sustaining these interactions are often left implicit. As (Schino & Aureli, 2010) note, reciprocity is frequently treated as a local response pattern, rather than a structured process involving memory and expectation. Yet reciprocal exchange, by definition, requires continuity. Without recognition or anticipation of future interaction, trust-like behavior cannot functionally emerge.
By defining these three mechanisms as simulateable behavioral primitives, we provide a bottom-up framework for modeling the emergence of economic exchange—one that avoids presupposing institutions, markets, or symbolic trust, and instead builds from capacities that are both empirically grounded and computationally implementable.
Figure 3: The agent integrates three input pathways via behaviorally grounded mechanisms—recognition, credence, and cost–return evaluation—to support dynamic decisions in social exchange contexts.
@@ 4.1 Individual Recognition
Definition 4.1 (Individual Recognition).
Individual recognition is the capacity to identify and distinguish specific social partners across time. It enables agents to track prior interactions with particular individuals, forming the basis for memory-based expectations about future behavior.
This capacity underlies the first and most basic source of reciprocal credence: without individual recognition, no information about past cooperation can be meaningfully retained or applied to guide future interaction.
Empirical Support.
@@ 4.2 Reciprocal Credence
To address this gap, we introduce the concept of reciprocal credence: a cognitively grounded, graded belief that another agent will respond to cooperation with cooperation. Unlike categorical trust, reciprocal credence is scalar, updateable, and context-sensitive. It is not a feeling, but a functional belief state—one that governs an agent’s willingness to initiate or sustain cooperative behavior.
Definition 4.2 (Reciprocal Credence).
Reciprocal credence is a scalar, updateable belief that another agent will respond to cooperation with cooperation. It is not a feeling, but a functional decision variable that governs an agent’s willingness to initiate or sustain cooperative behavior.
It arises from two temporally grounded types of inference:
  • Retrospective inference, based on memory of past prosocial behavior directed toward the self or others;
  • Prospective inference, based on contextual signals that indicate future reciprocity is likely—such as repeated interaction, social roles, or institutional incentives.
To make these sources operational in multi-agent systems, we distinguish three canonical informational pathways:
@@
    Direct interaction history, where an agent has experienced prior cooperative behavior, prosocial engagement, or affiliative signaling from a specific partner.
@@
    Reputational inference, based on observed or reported behavior toward others—supporting third-party inference of prosocial disposition.
@@
    Role-based or contextual expectation, where social roles, environmental incentives, or normative structures lead to default assumptions that cooperation will be returned, even in the absence of personal history.
These pathways differ in reliability, cognitive cost, and developmental trajectory—but together, they provide the informational grounding for reciprocal credence as a simulateable and biologically plausible substrate for reciprocal cooperation.
Empirical Support.
@@
The term “trust” is often applied to systems like ChatGPT or Google, but what is described is not social trust—it is functional reliability through repeated exposure and low failure rates. These systems do not engage in reciprocal reasoning or uphold social obligations.
This semantic conflation becomes especially problematic in high-stakes domains like finance and healthcare, where “trust” is better understood as a demand for exceptional systemic robustness—not social accountability.
In finance, phrases like “trust in banks” or “trust in money” reflect belief in institutional continuity, enforcement mechanisms, and liquidity guarantees—not interpersonal confidence. What matters is structural resilience against systemic collapse.
In healthcare, “trust in doctors” or “hospital systems” similarly refers to confidence in diagnostic accuracy, institutional safeguards, and low error rates. Patients rely on these systems not because of perceived benevolence, but because failures are rare and tightly bounded.
@@ 4.3 Cost–return sensitivity
Definition 4.3 (Cost–Return Sensitivity).
Cost–return sensitivity is the capacity to evaluate whether engaging in cooperative behavior yields a net benefit over time. Unlike categorical heuristics (e.g., always cooperate or always defect), this sensitivity enables agents to condition their participation on the payoff structure of ongoing interactions. It supports both the withdrawal from exploitative relationships and the reinforcement of beneficial ones, serving as a regulatory mechanism for reciprocal dynamics.
This capacity need not involve explicit calculation; approximate assessments—such as tracking recent imbalances in giving and receiving—are sufficient to support adaptive behavior.
Empirical Support.
@@
5 Simulation Pathways and Implementation
We define three simulateable behavioral primitives—individual recognition, reciprocal credence, and cost–return sensitivity—that can serve as the minimal cognitive capacities required for scalable reciprocal behavior in multi-agent systems. Each can be approximated through lightweight memory architectures and prompt-level reasoning, without relying on explicit reinforcement learning.
  • Individual recognition: Agents should be able to differentiate between social partners and retrieve information specific to each. This may reflect accumulated positive history, social familiarity, or bonded interaction patterns—not necessarily explicit naming.
  • Reciprocal credence: Agents should maintain an internal estimate of how likely a partner is to return cooperation. This estimate integrates: (1) prior prosocial behavior directed toward the agent or observed in others, and (2) contextual signals that make cooperative return likely. The value is asymmetric, updateable, and sensitive to both interaction and situation.
  • Cost–return sensitivity: Agents should track asymmetries in past exchanges, allowing them to adjust future cooperative investment. This mechanism supports dynamic calibration of effort and helps stabilize long-term reciprocity.
These primitives do not prescribe a specific architecture, but define necessary behavioral functions for any agent seeking to simulate socially grounded exchange. By shifting the modeling burden toward structured memory and belief inference, they offer a tractable and biologically plausible path for implementing reciprocal behavior in LLM-based systems.
@@ 6.1 Limitations of the Framework
Although we argue that modern exchange institutions build upon a behaviorally grounded behavioral substrate, we do not claim that such mechanisms alone are sufficient to explain the emergence of formal systems like currency, interest-bearing debt, or taxation. These likely require additional cultural, historical, and symbolic developments.
@@ 6.2 Theoretical Implications
Reciprocity as the Foundation of Social Structure.
The term cooperation is frequently used as a catch-all for prosocial behavior, but its vagueness makes it difficult to model or implement. Without a clear cognitive basis, many systems rely on artificial payoff matrices, fixed strategies, or exogenous incentives—mechanisms that simulate outcomes but not the behavioral processes behind them.
In contrast, reciprocity offers a cognitively grounded and behaviorally structured alternative. It is not merely repeated cooperation, but a mechanism based on individual recognition, interaction history, and cost–return regulation. These features make reciprocity tractable, simulateable, and aligned with real-world behavior.
More importantly, reciprocity underlies many core functions of modern society. From social favors and mentorships to business alliances and venture capital, cooperation often depends on asymmetric, memory-based, and deferred exchanges. Even scientific systems—such as peer review, citation, and referrals—operate through implicit reciprocity, not formal enforcement.
We argue that reciprocity—not abstract cooperation—is the minimal behavioral substrate of social structure. It enables agents to build conditional expectations, regulate ongoing interactions, and scale bilateral exchange into stable institutional patterns. Far from being a subset of cooperation, reciprocity is the mechanism that makes cooperation sustainable and socially extensible in the first place.
@@
To address this, we introduce reciprocal credence: a graded, behaviorally grounded expectation that cooperation will be returned. It emerges from recognition, interaction history, and cost–return sensitivity—not from moral obligation. In contrast, large-scale “trust” in systems like money or platforms is better understood as confidence in robustness, not social accountability. This reframing clarifies how reciprocal behavior scales without assuming institutional trust.
@@ 6.3.2 Simulating Institution Formation via Behavioral Primitives
With the rise of large language model (LLM)-based agents and embodied multi-agent simulators (e.g., VirtualHome (Puig et al., 2018), ThreeDWorld (Gan et al., 2020), AutoGen (Wu et al., 2023)), it is now possible to explore whether agents endowed with these primitives can spontaneously develop systems of reciprocal exchange, role specialization, symbolic debt, or proto-monetary behaviors. Such experiments would help identify the minimal substrate for scalable cooperation and provide a generative framework for building socially competent artificial agents.
@@ 7 Conclusion
Together, these primitives offer not just theoretical insight, but a tractable interface for embedding socially competent behavior in next-generation agent systems.